OpenStudy (yisselluna):

For the following system, use the second equation to make a substitution for x in the first equation. 3x + 2y = 7 x - y + 3 = 0 What is the resulting equation? 3x - y - 3 + 2y = 7 3(y - 3) + 2y = 7 3y - 3 + 2y = 7

OpenStudy (andrewishere):

Ok so simply make the second equation in terms of x, so we know what x equals.

OpenStudy (andrewishere):

This means we will have to move both negative y and positive 3 to the opposite side of the equation, as we want to keep x positive.

OpenStudy (andrewishere):

As we transfer both negative y and positive 3 to the other side, the signs change to their opposite. So... -y goes to t and +3 goes to -3.

OpenStudy (andrewishere):

Now we have the equation x=y-3

OpenStudy (andrewishere):

NOW we can substitute for x in the top equation. Sub in y-3 for x in the equation 3x+2y=7

OpenStudy (andrewishere):

Now we have 3(y-3) + 2y = 7

OpenStudy (andrewishere):

Use the distributive property for 3(y-3). This gives us 3y-9. Now add this to the rest of the equation. 3y-9+2y = 7

OpenStudy (andrewishere):

Oh I'm sorry, we don't have to go further than 3(y-3)+2y = 7

OpenStudy (andrewishere):

DId this help?

OpenStudy (yisselluna):

yes thanks!

OpenStudy (andrewishere):

No problem